# Neurodynamics-Driven Coupled Neural P Systems for Multi-Focus Image Fusion

### v2

@@ 5.1 Experimental configurations
Dataset. To fully evaluate the effectiveness of NDCNPFuse, we conducted extensive experiments covering four datasets: Lytro [21], MFFW [34], MFI-WHU [37], and Real-MFF [38]. The Lytro dataset includes 20 pairs of color images. The MFI-WHU dataset consists of 120 pairs of artificially synthesized color images. The MFFW dataset comprises 13 pairs of color images with defocus spread effects (DSE). The Real-MFF dataset contains 710 pairs of real-world MFIF images.
@@
Implementation details. ND-CNPFuse runs on an Intel(R) Core(TM) i5-13400 CPU, and DL methods run on an Nvidia A100 GPU. We set the coupling radius $r$ to 16 and iterations $t$ to 110. Please see Sec. 5.3 for detailed parameter settings.
@@ 5.2 Performance comparison
We further visualize the difference images to compare fusion results, as shown in Fig. 6. Most methods show visible residuals, while PADCDTNP introduces boundary artifacts (see white boxes). In contrast, ND-CNPFuse fully preserves the focusing information from the source images, achieving more accurate boundaries.
Quantitative comparison. Tab. 1 reports the quantitative results of all methods on four datasets. ND-CNPFuse ranks first in five of six evaluation metrics across four datasets. It falls short of the best PSNR on MFFW as its spike count-based clarity measure may inadequately capture low-contrast edges, slightly impacting edge preservation. Nevertheless, it still ranks second. This indicates that our method has good generalizability. We also acknowledge that the numerical improvement over the second-best method is not significant. However, the contribution of our work lies not only in performance gains but also in offering a novel research perspective for the MFIF field.
@@ 5.3 Parameter sensitivity analysis
Impact of coupling radius. We test coupling radius values of 2, 4, 8, 16, and 32. As shown in Fig. 77(a), $Q_{percep}$ increases with radius and peaks at 16, then declines due to loss of detail and gradient information. Thus, $r$ is set to 16.
@@ 5.4 Ablation study
Neurodynamic analysis. We conducted ablation studies on the Lytro dataset to validate the effectiveness of neurodynamic analysis in improving model performance. Specifically, we compare the baseline CNP system and the neurodynamics-driven CNP system across six metrics, as summarized in Tab. 3. We can find that the neurodynamics-driven CNP system achieves significant improvements over the baseline across six multidimensional image quality metrics, with improvements of 2.01%, 17.29%, 1.55%, 5.01%, 1.44%, and 7.29%. These findings confirm the effectiveness of neurodynamic analysis.
SML module. We further validated the impact of the SML module on overall architecture performance. According to the quantitative analysis results in Tab. 4, after removing the SML module, evaluation metrics showed only a slight downward trend (e.g., PSNR decreased by 0.007 dB, and SSIM decreased by 0.002). This confirms the generalizability of the proposed method. Of course, by applying SML, the quality of fusion can be further enhanced.
@@ Acknowledgement
This work was supported by National Key Research and Development Program of China (No. 2022YFC3303600), the National Natural Science Foundation of China (No. 62137002, 62293553, 62293554, 62437002, 62477036, and 62477037), the “LENOVO-XJTU” Intelligent Industry Joint Laboratory Project, the Fundamental Research Funds for the Central Universities (No. xzy022024056), the Shaanxi Provincial Social Science Foundation Project (No. 2024P041), the Natural Science Basic Research Program of Shaanxi (No. 2023-JC-YB-593), the Youth Innovation Team of Shaanxi Universities “Multi-modal Data Mining and Fusion”, Shaanxi Undergraduate and Higher Education Teaching Reform Research Program (No. 23BY195), Xi’an Jiaotong University City College Research Project (No. 2024Y01), and China Knowledge Centre for Engineering Sciences and Technology.

### v3

@@ 5.1 Experimental configurations
Dataset. To fully evaluate the effectiveness of NDCNPFuse, we conducted extensive experiments covering four datasets: Lytro [25], MFFW [38], MFI-WHU [40], and Real-MFF [43]. The Lytro includes 20 pairs of color images. The MFFW comprises 13 pairs of color images with defocus spread effects (DSE). The MFI-WHU consists of 120 pairs of artificially synthesized color images. The Real-MFF contains 710 pairs of real-world MFIF images.
@@
Implementation details. ND-CNPFuse runs on an Intel(R) Core(TM) i5-13400 CPU, and DL methods run on an Nvidia A100 GPU. We set the coupling radius $r$ to 16 and the number of iterations $t$ to 110, as detailed in Sec. 5.3.
@@ 5.2 Performance comparison
We further visualize the difference images to compare fusion results, as shown in Fig. 6. Most methods show visible residuals, while PADCDTNP introduces boundary artifacts. In contrast, ND-CNPFuse fully preserves the focusing information from the source images. We also evaluate downstream salient object detection tasks in Appendix C.2.
Quantitative comparison. Tab. 1 reports the quantitative results of all methods on four datasets. ND-CNPFuse ranks first in five of six evaluation metrics across four datasets. It falls short of the best SSIM on MFFW as its spike count-based clarity measure may inadequately capture low-contrast edges, slightly impacting edge preservation. Nevertheless, it still ranks second. We acknowledge that the numerical improvement over the second-best method is not significant. However, the contribution of our work lies not only in performance gains but also in offering a novel research perspective for the MFIF field. We also include the classic spiking neurons LIF model [5] for comparison, with detailed results provided in Appendix C.3.
@@
Energy consumption. Following the energy calculation in [9], the energy consumption for ND-CNPFuse on the Lytro dataset ($520\times 520$) is $1.12\times 10^{-5}$J per image pair.
@@ 5.3 Parameter sensitivity analysis
Impact of coupling radius. We test coupling radius values of 0, 2, 4, 8, 16, and 32. As shown in Fig. 77(a), $Q_{percep}$ increases with radius and peaks at 16, then declines due to loss of detail and gradient information. Thus, $r$ is set to 16.
@@ 5.4 Ablation study
Neurodynamic analysis. We conducted ablation studies on the Lytro dataset to validate the effectiveness of neurodynamic analysis in improving model performance. Specifically, we compare the baseline CNP system and the neurodynamics-driven CNP system across six metrics, as summarized in Tab. 3. We can observe that the neurodynamics-driven CNP system achieves significant improvements over the baseline across six multidimensional image quality metrics, with improvements of 2.01%, 17.29%, 1.55%, 5.01%, 1.44%, and 7.29%. These findings confirm the effectiveness of neurodynamic analysis.
SML module. We further validated the impact of the SML module on overall architecture performance. According to the quantitative results in Tab. 4, after removing the SML module, evaluation metrics showed only a slight decline. This confirms the generalizability of the proposed method. Of course, by applying SML, the quality of fusion can be further enhanced. Moreover, SML is beneficial for handling noisy inputs. More details can be found in Appendix C.4.
@@ Acknowledgement
This work was supported by National Key Research and Development Program of China (No. 2022YFC3303600), the National Natural Science Foundation of China (No. 62137002, 62293553, 62293554, 62437002, 62477036, and 62477037), the “LENOVO-XJTU” Intelligent Industry Joint Laboratory Project, the Fundamental Research Funds for the Central Universities (No. xzy022025037), the Shaanxi Provincial Social Science Foundation Project (No. 2024P041), the Natural Science Basic Research Program of Shaanxi (No. 2023-JC-YB-593), the Youth Innovation Team of Shaanxi Universities “Multi-modal Data Mining and Fusion”, Shaanxi Undergraduate and Higher Education Teaching Reform Research Program (No. 23BY195), Xi’an Jiaotong University City College Research Project (No. 2024Y01), and China Knowledge Centre for Engineering Sciences and Technology.